What is the maximum distance from which a sailor on top of the
mast of ship 1, at 34 m above the ocean’s surface, can see another sailor on
top of the mast of ship 2, at 26 m above the ocean’s surface?
1
Expert's answer
2021-09-28T11:13:17-0400

The vertical distance between the two sailors is


"h=34-26=8\text{ m}."

The distance in km is


"d\approx3.57\sqrt h=10\text{ km}."